# Efficiently Serving LLM Reasoning Programs with Certaindex

### v1

@@ 6.1 Experimental Setup
Batch processing and metrics. For offline workloads, we measure tokens-to-accuracy, the total number of generated tokens needed to achieve specific accuracy levels across reasoning queries. This metric reflects the accuracy-cost trade-off, critical for end users as LLM platforms charge based on token usage. Table 2 details the offline experimental settings. Each problem is allocated a maximum computation budget (parameterized by a Resource Cap) defining the sampling branch limit for SC, with limit for Rebase, and iteration limit for MCTS. Resource caps correspond to different points on the same line in Fig. 8. The resource cap in our experiments may correspond to the reasoning_effort configuration parameter in o1 [37], as a higher resource cap means a higher accuracy. While scheduling methods may terminate programs early, they cannot exceed these caps.
@@
Online serving and metrics. For online workloads, we simulate query arrivals using dataset queries and assign deadlines to each query. System performance is evaluated by deadline attainment, the percentage of queries completed within their deadlines. We vary arrival rates and SLO scales to assess system performance under different conditions. Tab. 3 outlines the online experiment settings. Request arrivals follow a Poisson process with varying rates. Deadlines are difficulty-aware, determined using a simple policy: oracle difficulty for each query is estimated through sufficient number of trial runs for each algorithm-dataset combination, classifying queries as always correct (difficulty factor $=1$), always incorrect ($=3$), or variable ($=2$). Deadlines are calculated as the product of the SLO scale, the difficulty factor, and a base deadline, accounting for the additional time required for more challenging problems.
6.2 Batch Processing
@@
Baselines. For offline evaluation, we compare Dynasor against a modified SGLang intra-program scheduler using the following policies: (1) baseline-even. It allocates resources uniformly across all reasoning programs using the resource cap settings in Tab.2. Different cap values result in different accuracy levels, as larger caps allow more computation for all problems. Specifically, for each algorithm, all queries receive identical resources: SC uses the same number of branches, Rebase uses the same width, and MCTS uses the same number of search iterations, as specified in Tab. 2. This baseline reflects the behavior of reasoning programs in current systems: user may allocate equal resources on all problems. (2) baseline-length. This policy uses the cumulative token count generated at a specific step (specified as Detect@knob in Table 1) as the program’s ending signal. The detection step matches the one used in the certaindex-based approach. The scheduler continues resource allocation if the historical token count over a specified number of steps exceeds a predefined threshold, indicating the problem’s difficulty requires longer reasoning chains with more tokens. Output token length is commonly used as a scheduling indicator in existing LLM serving systems [13, 23, 50]. It also correlates strongly with the compute requirements (knob size) as we have discussed in §3.3.
(Algo., Dataset, LLM) Reward Model # Samples Resource Cap
(SC, LCB, Llama3.1 8B) / 400 5,10,15,20,25,30
(SC, GSM8K, Llama3.1 8B) / 1000 5,10,15,20,25,30
(MCTS, ASDiv, Llama2 7B) Skywork 7B 300 3,7,10,15,20
(MCTS, GSM8K, Llama2 7B) Skywork 7B 300 3,7,10,15,20
(Rebase, MATH, Llemma 7B) Llemma 34b 500 16,32,64,128
(Rebase, GSM8K, Llemma 34B) Llemma 34b 500 16,32,64,128
Fig. 8 compares the token-to-accuracy metric on different workloads. Our method consistently outperforms all baselines across different workloads by reducing token usage by 9-48% compared to baseline-even and 9-52% compared to baseline-length without sacrificing accuracy. On SC-LiveCodeBench, Dynasor can reduce token usage by 9.5% compared to baseline-even, and reduce up to 21% token usage compared to baseline-length, without accuracy dropping. On SC-GSM8K, Dynasor achieves up to 47.8% and 47% token savings vs. baseline-length and baseline-even, respectively. On MCTS, our method shows 9% and 11.8% fewer tokens usage compared to baseline-length, 13% and 13.1% fewer tokens usage compared to baseline-even on ASDiv and GSM8k respectively. For Rebase, we obtain 52.2% and 11.3% token reduction compared to baseline-length and baseline-even on Math, and 43.5% and 48% token reduction compared to baseline-length and baseline-even respectively on GSM8K.
These gains primarily come from the intra-program scheduling algorithm, which accurately identifies high-certaindex programs and terminate them early without accuracy loss, as explained in §3 and calibrated in §5.2.2. This early termination strategy significantly reduces resource consumption by eliminating unnecessary sampling compared with baseline-even. In contrast, baseline-length leads to accuracy degradation even with less aggressive compute pruning, highlighting the effectiveness of certaindex-based resource allocation.
Table 3: Online workload Configurations
Algorithm Dataset LLM Reward Model Base Deadline (s)
SC MATH Llama3.1 8B / 240
MCTS ASDiv Llama2 7B Skywork 7B 60
Rebase GSM8K Llemma 34B Llemma 34b 300
6.3 Online Serving
@@
$\bullet$ SGLang [57]. SGLang represents a strong baseline as it incorporates key optimizations in LLM serving: its longest-prefix matching (LPM) algorithm efficiently batches requests with common prefixes, while prefix caching reduces redundant computation.
@@
$\bullet$ ParrotServe [29]. ParrotServe uses gang scheduling (App-FIFO) to prioritize requests in the same program. We implemented its scheduler on top of SGLang for fair comaprison. By grouping requests from the same program together, it minimizes context switches and maximizes KV cache reuse across batches. This baseline applies to gang scheduling at the program level.
Figure 9: Evaluation on 3 online workloads on Dynasor against baselines. Rows from top to bottom: (a) Program arrival rate vs SLO attainment, (b) SLO scale vs SLO attainment, (c) Accuracy vs SLO Attainment.
@@
Fig. 9 presents three key performance trade-offs: (a) Program arrival rate, (b) SLO scale, and (c) Accuracy vs. SLO attainment. For experiments (a) and (b), we use fixed resource cap settings (SC: 20, MCTS: 15, Rebase: 128) to compare Dynasor with baselines, which lack adaptive compute scheduling for LLM reasoning programs. We measure SLO attainment while ensuring accuracy is maintained. Our method employs the same resource allocation policy validated in §6.2 and §A, which preserves accuracy. In experiment (c), we explore the relationship between achievable accuracy and SLO attainment by varying resource cap settings.
Rate vs. SLO attainment. Our system achieves much higher sustainable request rates under P90 deadline attainment: 1.63.3×1.6-3.3$\times$1.6 - 3.3 × and 1.63.2×1.6-3.2$\times$1.6 - 3.2 × compared to SGLang and Parrot respectively. These gains stem from two key factors. First, our intra-program scheduler identifies and early-terminates programs with high confidence, freeing resources for pending requests. Second, our inter-program scheduler prioritizes requests from the same program and enables program-level schedule (e.g., program-level SJF), increasing batch turnover rate. While our approach requires two rounds of execution with minimal knobs to detect certaindex, potentially impacting parallelism, it still delivers substantial performance gains: approximately 1.6×1.6$\times$1.6 × for SC and $\times3$ for Rebase, even in these highly parallel workloads where requests could be issued all at once. As shown in Fig. 9, SGLang and Parrot achieve similar performance on MCTS due to its inherently lower parallelism, which limits the impact of scheduling optimizations.
SLO scale vs. SLO attainment. Fig. 9(b) demonstrates how different systems maintain deadline attainment under a fixed request rate (8, 16, 2 programs/s for SC, MCTS, Rebase). At these rates, Dynasor allows tighter SLO scales: 1.34.7×1.3-4.7$\times$1.3 - 4.7 × tighter than SGLang, and 1.73.3×1.7-3.3$\times$1.7 - 3.3 × tighter than Parrot.
Accuracy vs SLO attainment. We measure system performance under fixed request rates (8, 16, 1 programs/s for SC, MCTS, Rebase) across different pre-configured resource cap settings (as in Tab. 2), where each resource cap determines the maximum computational budget and consequently affects serving accuracy. Fig. 9 (c) shows the trade-off between accuracy and SLO attainment across these configurations. On all three workloads, Dynasor shows 0.7% - 2% accuracy improvement compared to SGLang and Parrot under the same SLO attainment. This improvement stems from our ability to redistribute compute between simple and hard queries, enabling us to solve more queries while maintain similar SLO attainment that baselines could not match unless given much more compute.
Throughput. Our system shows equal average throughput (token per second) in all online settings to baselines. This is because that under the given request rates, all workloads saturate the GPU memory and making the system memory bound. This also validates the introduced scheduler has no observable overhead.
@@
6.4 Fine-grained Resource Allocation
We compare simple thresholding against fine-grained resource allocation in this section. §6.2 and §6.3 adopt a simple static threshold mechanism: extract certaindex at a fixed step and decide if to terminate program based of the value of certaindex. While effective, there is room for optimization through more sophisticated scheduling strategies. To explore this, we conducted experiments in offline batch processing using the setting (SC, MATH) and (MCTS, ASDiv), both with a maximum step limit of 20 as per-query resource cap, to evaluate more complex allocation strategies that enable early termination while maintaining accuracy. The results, measured in token savings, are presented in Tab. 4. For the static threshold, we collected certaindex and apply the threshold in Tab. 1. For all following methods, we calibrate the mapping between certaindex and compute requirements across 10 different runs, ensuring no accuracy drop for each method in all runs. We then evaluate different methods’ token savings by measuring compute requirements for each problem in the dataset using their empirical certaindex.
We first examine an Initial Step Curve Fitting approach, which fits a skyline curve (exemplified by the green lines in Fig. 3) using certaindex values collected at the same steps as the static threshold. Although this method enables finer-grained resource allocation for varying certaindex values, it relied on early step’s certaindex values, which may be inaccruate as reasoning progresses, resulting in marginal improvements (less than 1% compute savings) compared to simple static thresholding.
6.5 Ablation Studies
We ablate the effectiveness of two core proposed techniques: resource scheduler and gang scheduling algorithm.
Components Contributions. The speedup of the system comes from two major sources: (1) intra-program scheduling with dynamic resource allocation and (2) inter-program scheduling leveraging program-level locality. Figures 10 and 11 show the contribution of different components using SC workloads on GSM8K and MATH datasets, measuring mean latency and maximum sustainable rate under the same P90 attainment, respectively. Using LPM (SGLang) as the baseline, we analyze three factors: gang scheduling, SJF, and certaindex-based resource management. Gang scheduling improves latency and throughput in both uniform and certaindex-aware resource allocation by ensuring the requests in the same program are scheduled together, reducing resource fragmentation and context switching overhead. For GSM8K workload, certaindex-based resource management dominates the improvements, reducing mean latency by up to 60% through token savings of up to 50%, while Gang scheduling and SJF contributions are more modest. For MATH workload, certaindex-aware allocation achieves a 1.2x peak rate through early termination of low-value computations (as only 10% tokens are saved on MATH), while SJF provides the most significant gain with a 1.9x peak rate by prioritizing shorter jobs and reducing head-of-line blocking. These results validate the effectiveness of our two-level scheduling approach: intra-program optimization through certaindex-aware allocation, combined with inter-program optimization through gang scheduling.
Figure 10: Performance improvement breakdown in online self-consistency (GSM8K) workload: Impact of gang scheduling, certaindex-based resource allocation, and SJF on mean latency given a fixed request rate of 16 pps.
Choosing Different Thresholds. We show that certaindex threshold selection is vital. Fig. 12 demonstrates the impact of different certaindex thresholds. For SC on GSM8k, we selected an entropy ($\mathcal{H}$) threshold of 0.5, as higher thresholds led to overly aggressive termination and accuracy degradation. Similarly, for MCTS on ASDiv, we chose a score ($\mathcal{R}$) threshold of 0.4, as lower thresholds (<0.4) decreased accuracy while higher thresholds (>0.4) increased token consumption without proportional accuracy gains. These results highlight the importance of careful threshold policy calibration in balancing compute efficiency and accuracy as we discussed in §5.2.2.
Comparison to other signals. We further compare our certaindex-based approach with two alternative signals that may guide inference-time compute allocation: length-based [13] and LLM activation-based methods [8]. The first has been discussed in §3.3 and §6.2 and is a simple and popular scheduling indicator [23, 13] in existing systems. We carefully tune it with varying lengths. The activation-based method [8], which uses a neural network predictor trained on LLM activation with workload-specific data, to predict queries’ resource requirement. Conceptually, it shall capture similar or more signals than what certaindex provides. In Figure 13, our certaindex-based method consistently outperforms both across workloads, even achieving higher accuracy with lower number of tokens. The length-based method, which terminates based on generation sequence length, either sacrifices accuracy with higher cutoffs or shows minimal token savings with lower cutoffs. The activation-based method shows accuracy degradation in SC tasks and poor token efficiency in MCTS – model-based resource predictor is non-trivial to train and suffer from query distribution shift. These comparisons validate the simplificy and effectiveness of our certaindex-based resource scheduling and align with our findings in §3.3 that certaindex is more powerful in guide reasoning program resource allocation.
Figure 13: Performance comparison with LLM activation-based predictor and output length based scheduling.
6.6 Fairness Analysis
We evaluate finish-time fairness (§5.3.1) using the ratio of latency to baseline token count, where baseline tokens represent the amount needed when all programs run to their maximum resource cap - this is equivalent to the actual tokens used in systems without our optimization, and serves as the reference point for systems using our optimization. Fig. 14 shows the comparison on certaindex-based resource scheduling, gang scheduling, and SJF on finish-time fairness on MATH using SC with a rate of 8 pps using Llama3.1 8B. Gang scheduling shows consistent fairness improvement compared to non-gang scheduling. This shows that prioritizing existing programs can improve finish-time fairness. Certaindex-based resource allocation also consistently improve finish-time fairness due to resource cutting by the intra-program scheduler. Adding SJF shows fairness improvement at the later 50% fraction of the job compared to without SJF, and at the later 65% fraction of job compared to LPM. In all case, SJF shows the fairness metric no worse than even resource allocation. Note that despite improvements in finish-time fairness, starvation may still occur. We provide options to mitigate this issue, as discussed in §5.3.1.
Since GPU memory capacity is an important resource consideration in today’s LLM serving systems, we also study the how varying GPU memory affects the SLO attainment, and defer the experiments to Appendix B.
6.7 Case Study: Serving o1-like LLMs
We conduct a case study to evaluate Dynasor’s effectiveness in serving models with reasoning behaviors similar to OpenAI’s lastest o1. Since o1 is proprietary, at the time of submission, we use the closest open source LLM – Qwen-QWQ [47], which is trained with internalized CoT and shows comparable performance to o1. We find QWQ generates lengthy single CoT paths with continuous reflection. We evaluate the model’s tokens-to-accuracy performance by varying token generation limits across 500 GSM8K problems. The original model, which only produces answers after completing entire CoT paths, showed poor tokens-to-accuracy performance under lower token limits (green curve in Fig. 15). This limitation arises because the model frequently fails to reach conclusions within constrained budgets.
To improve it, we let QWQ generate 32 tokens for each problem per iteration and inject a “Final Answer” prompt to instruct the model to elicit answers at the end of every iteration (QWQ Guided). This approach reduces token consumption by 30% while maintaining the same accuracy. Finally, we monitors the certaindex (specifically, $\tilde{\mathcal{H}}$) of answers from different iterations and terminates generation when the certaindex exceeds a confidence threshold (QWQ Cut). QWQ Cut achieves an additional 14% reduction in token consumption without accuracy loss, demonstrating the effectiveness of our approach for efficiently serving latest models where reasoning behaviors are internalized.
Figure 14: Finish-Time Fairness.
Figure 15: Serving Qwen-QWQ using Dynasor.
7 Related Work
We discuss more related work besides those in §2.3.
Goodput-Optimized Serving Systems. Optimizing goodput has become a recent focus in machine learning systems. Symphony [6], Nexus [43], Clockwork [14], and Shepherd [56] adopt eager scheduling with preemption to improve the goodput of serving traditional deep neural networks. Conceptually, the closest to our work is Pollux [40], which optimizes a goodput objective that balances the system throughput and statistical efficiency at training. Dynasor can be also seen as a goodput-optimized system, but focused at inference. It dynamically manages inference-time compute for each reasoning query to minimize latency and maximize throughput while preserving overall accuracy.
Adaptive resource allocation for LLM queries. Recent studies [45, 26, 36, 8, 32] explore adaptive compute allocation for LLM inference based on query difficulty. RouteLLM [36] routes queries between multiple LLMs to reduce cost, but it requires prior knowledge of traffic patterns. Capability-Aware Self-Evaluation [32] estimates the likelihood of failing to generate preferred responses. Damani et al. [8] predicts the reward distribution for inputs under a compute budget, allocating more resources to promising queries. However, this approach depends on trained predictors that are hard to obtain and vulnerable to query distribution shifts. Sun et al. [45] examine the relationship between reward model outputs and inference-time resource allocation, providing an alternative measurement of certaindex. Specific to SC, Li et al. [26] show that many SC programs can be early exited using a sliding window detector without compromising accuracy, echoing our findings. Building on these ideas, we propose certaindex, a simple, plug-and-play indicator that generalizes across diverse reasoning algorithms and datasets.
8 Conclusion
We introduced Dynasor for adaptively orchestrating inference-time compute for LLM reasoning programs. By leveraging certaindex as a proxy for reasoning progress, Dynasor optimizes compute allocation across diverse reasoning queries. Our evaluations demonstrate significant performance gains over existing systems. We hope Dynasor sets a new bar for serving the emerging LLM reasoning programs.
Appendix A Token-to-accuracy Performance of Dynasor on MATH using SC

### v2

@@ Appendix F Detailed Setups for Offline and Online Experiments
Compute. All experiments run on a GPU cluster (Runpod) equipped with A100 (80GB) GPUs. GPU usage varies by method: Rebase and MCTS use two GPUs to separately serve their LLM and reward models, while SC requires only one GPU for request processing. For final answer selection, SC uses simple majority voting, Rebase applies weighted majority voting, and MCTS selects the best path using a reward model via best-of-n.
General settings. Tables 1 and 2 summarize our offline and online evaluation workloads, respectively. Table 3 lists the certaindex hyperparameters for each workload, including the thresholds at the detection step (detect @knob). A program terminates at the detection step if its certaindex values meet all threshold conditions.
F.1 Batch Inference Setup: CoT
@@
Metrics. For offline workloads, we measure tokens-to-accuracy, the total number of generated tokens needed to achieve specific accuracy levels across reasoning queries. This metric reflects the accuracy-cost trade-off, critical for end users as LLM platforms charge based on token usage. Table 1 details the offline experimental settings for SC/MCTS/REBASE. Each problem is allocated a maximum computation budget (parameterized by a Resource Cap) defining the sampling limit for SC and iteration limit for MCTS. Resource caps correspond to different points in Fig. 7. While scheduling methods may terminate programs early, they cannot exceed these caps. Query deadlines are calculated as the product of three factors: the SLO scale, the query’s difficulty factor, and a base deadline, reflecting the additional time required for more challenging problems.
We evaluate the Certaindex-based early termination method (§2.1) against baseline uniform token allocation across multiple scales of distilled DeepSeek models (7B, 14B, and 32B) [14] on mathematical reasoning benchmarks AIME24 and AMC23 [20], and MATH500 [39]. Unlike the baselines that uniformly increases token budgets, our appraoch early terminates by monitoring Certaindex at various intervals ($T$ = 32, 64, 128. 256, and 320). All requests are given max token budget of 16K. For each interval, we vary the early termination parameter $N$ (the required number of consecutive consistent answers), generating different points along each line. For fair comparison, appropriate accuracy thresholds were calibrated to model scale - with 32B models evaluated against stricter thresholds above QwQ [34] levels and reduced thresholds for smaller models - while setting higher targets for simpler tasks where greater accuracy is achievable.
F.2 Batch Inference Setup: SC, MCTS, Rebase
@@
Metrics. For online workloads, we simulate query arrivals using dataset queries and assign deadlines to each query. System performance is evaluated by P90 deadline attainment, the percentage of queries completed within their deadlines. We vary arrival rates and SLO scales to assess system performance under different conditions. Tab. 2 outlines the online experiment settings. Request arrivals follow a Poisson process with varying rates. Deadlines are difficulty-aware, determined using a simple policy: oracle difficulty for each query is estimated through extensive trial runs (> 100) for each algorithm-dataset combination, classifying queries as always correct (difficulty factor $=1$), always incorrect ($=3$), or variable ($=2$). Deadlines are calculated as the product of the SLO scale, the difficulty factor, and a base deadline, accounting for the additional time required for more challenging problems.
Baselines. For offline evaluation, we compare Dynasor against a modified SGLang intra-program scheduler using the following policies:
@@
$\bullet$ (1) baseline-even, allocates resources uniformly across all reasoning programs using the resource cap settings in Tab.1. Different cap values result in different accuracy levels, as larger caps allow more computation for all problems. Specifically, for each algorithm, all queries receive identical resources: SC uses the same number of branches, Rebase uses the same width, and MCTS uses the same number of search iterations, as specified in Tab.1. This baseline reflects the behavior of reasoning programs in current systems: user may allocate equal resources on all problems.
$\bullet$ (2) baseline-length. This policy uses the cumulative token count generated at a specific step (specified as Detect@knob in Table 3) as the program’s progress signal. The detection step matches the one used in the certaindex-based approach. The scheduler either continues allocating resources or terminates the program based on a predefined token-length threshold.
Output token length is commonly used as a scheduling indicator in existing LLM serving systems [41, 46, 61]. It also correlates strongly with the compute requirements (knob size) as we have discussed in §B.3.
F.3 Online Inference Detailed Setup
@@
$\bullet$ SGLang [35]. SGLang represents a strong baseline as it incorporates key optimizations in LLM serving: its longest-prefix matching (LPM) algorithm efficiently batches requests with common prefixes, while prefix caching reduces redundant computation. We enabled these optimizations and tuned for stable performance with 70% memory utilization.
@@
$\bullet$ Parrot [40]. Parrot uses gang scheduling (App-FIFO) to prioritize requests in the same program. We implemented its scheduler on top of SGLang for fair comaprison. By grouping requests from the same program together, it minimizes context switches and maximizes KV cache reuse across batches.
@@
For experiments (a) and (b), we use fixed resource cap settings (SC: 20, MCTS: 15, Rebase: 128) to compare Dynasor with baselines, which lack adaptive compute scheduling for LLM reasoning programs. We measure SLO attainment while ensuring accuracy is maintained. Our method employs the same resource allocation policy validated in §4.1 and §H, which preserves accuracy. In experiment (c), we explore the relationship between achievable accuracy and SLO attainment by varying resource cap settings.
Appendix G Other Ablation Studies
G.1 Effectiveness of CoT in Deepseek-R1
To validate scalability, we extended our experiments to the larger DeepSeek-R1 model on AIME and AMC datasets (Figure 17). The results align with our findings from smaller distill models, demonstrating consistent efficiency gains: DeepSeek-R1 achieves 12% token savings on AIME problems and 24% on AMC problems while maintaining baseline accuracy levels.
Figure 17: Apply Dynasor on DeepSeek-R1
G.2 Ablation: Scheduling Component Contributions
Figure 18: Performance improvement breakdown in online SC (GSM8K): Impact of gang scheduling, certaindex-based allocation, and SJF on mean latency with fixed request rate (rps = 16).
Figures 19 and 19 show the contribution of different components using SC on GSM8K and MATH datasets, measuring mean latency and maximum sustainable rate under the same P90 attainment. Using LPM (SGLang) as the baseline, we analyze three factors: gang scheduling, SJF, and certaindex-based resource management. Gang scheduling improves latency and throughput in both uniform and certaindex-aware resource allocation by ensuring the requests in the same program are scheduled together, reducing resource fragmentation and context switching overhead. For GSM8K workload, certaindex-based resource management dominates the improvements, reducing mean latency by up to 60% through token savings of up to 50%, while Gang scheduling and SJF contributions are more modest. For MATH workload, certaindex-aware allocation achieves a 1.2x peak rate through early termination of low-value computations (as only 10% tokens are saved on MATH), while SJF provides the most significant gain with a 1.9x peak rate by prioritizing shorter jobs and reducing head-of-line blocking. These results validate our two-level scheduling approach: intra-program optimization through gang scheduling and certaindex-aware allocation, combined with inter-program optimization through SJF.
G.3 Fairness Analysis
We apply finish-time fairness (§D.3.1) as the fairness metric and use latency / number of tokens as a proxy. Figure 20 shows the comparison on certaindex-based resource scheduling, gang scheduling, and SJF on finish-time fairness on MATH using SC with a rate of 8 pps using Llama3.1 8B. Gang scheduling shows consistent fairness improvement compared to non-gang scheduling. This shows that prioritizing existing programs can improve finish-time fairness. Certaindex-based resource allocation also consistently improve finish-time fairness due to resource cutting by the intra-program scheduler. Adding SJF shows fairness improvement at the later 50% fraction of the job compared to without SJF, and at the later 35% fraction of job compared to LPM. In all case, SJF shows the fairness metric no worse than even resource allocation.
Figure 20: Finish-Time Fairness.
G.4 Fine-grained Resource Allocation
We compare simple thresholding against fine-grained resource allocation in this section. §4.1 and §4.2 adopt a simple static threshold mechanism: extract certaindex at a fixed step and decide if to terminate program based of the value of certaindex. While effective, there is room for optimization through more sophisticated scheduling strategies. To explore this, we conducted experiments in offline batch processing using the setting (SC, MATH) and (MCTS, ASDiv), both with a maximum step limit of 20 as per-query resource cap, to evaluate more complex allocation strategies that enable early termination while maintaining accuracy. The results, measured in token savings, are presented in Tab. 4. For the static threshold, we collected certaindex and apply the threshold in Tab. 3.
We first examine an Initial Step Curve Fitting approach, which fits a skyline curve (illustrated by the green lines in Fig. 12) using certaindex values collected at the same steps as the static threshold. Although this method enables finer-grained resource allocation for varying certaindex values, it relied on early certaindex values, which maybe inaccruate as reasoning progresses, resulting in marginal improvements (less than 1% compute savings) compared to simple static thresholding.
@@ G.4 Fine-grained Resource Allocation
Appendix H Token-to-accuracy Performance of Dynasor on MATH using SC